Circular No. 7111

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 D1 The LONEOS program reports the discovery of an apparent comet.

The following observations are available:

1999 UT α_{2000} δ_{2000} m_{1} Observer
Feb. 20.40541 12^{h}15^{m}00.^{s}34 +3^{o}20′24′′.3 18.4 Hermann
20.42580 12 14 59.55 +3 19 50.7  _{”}
20.44621 12 14 58.73 +3 19 16.0  _{”}
20.94853 12 14 40.52 +3 05 28.0  Pravec
20.95084 12 14 40.40 +3 05 24.8  _{”}
20.95317 12 14 40.29 +3 05 19.8  _{”}
20.95551 12 14 40.17 +3 05 16.4  _{”}
20.95784 12 14 40.03 +3 05 13.2  _{”}
20.96249 12 14 39.90 +3 05 04.8  _{”}

S. M. Hermann (Lowell Observatory). 0.59-m LONEOS Schmidt telescope

+ CCD. Measurer B. W. Koehn. P. Pravec (Ondřejov). 0.65-m f/3.6 reflector + CCD. Comet has a broad

tail 15′′ long in p.a. 290^{o}.

GRB 990217 L. Piro, Istituto Astrofisica Spaziale (IAS), Rome; L. A. Antonelli, Os-

servatorio Astronomico, Rome; G. D’Andreta and V. Torroni, BeppoSAX Scientific Operation Center (SOC), Rome; M. Dadina, IAS and SOC, Rome; G. Gennaro, BeppoSAX Operation Control Center, Telespazio, Rome; and M. R. Daniele, BeppoSAX Science Data Center, Rome, report: “The BeppoSAX Wide Field Camera (WFC) error box of GRB 990217 (IAUC 7110) was observed starting on Feb. 17.49 UT (6 hr after the γ-ray trigger), using the Narrow Field Instruments (NFI) of BeppoSAX. After a preliminary analysis of the first 15 hr of observation, no source is detected, with a 2σ upper limit of ∼ 1 × 10^{−13} erg cm^{−2} s^{−1} in the range 2–10 keV. The slope of the decay power law connecting the WFC and NFI should then be steeper than ∼ −1.6 (i.e., more steep than in all the x-ray afterglows observed so far with BeppoSAX).”

NOVA MUSCAE 1998 Visual magnitude estimates by A. Pearce, Dowerin, W. Australia: Jan.

17.81 UT, 12.4; 18.81, 12.5; 20.80, 13.3; Feb. 17.74, 14.2.

1999 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green